# DietGlance: Dietary Monitoring and Personalized Analysis at a Glance with Knowledge-Empowered AI Assistant

### v2

@@ 5.3.3 Results
5.3.4 Post-Analysis Discussion with Experts
We held a panel discussion with the experts to analyze the results with two themes: (1) reasons for low performance in specific nutrients and (2) key nutrients for dietary analysis and suggestions.
@@ 5.3.4 Post-Analysis Discussion with Experts
(2) The experts valued the system’s overall performance for daily dietary tracking, since the important nutrients, such as the macronutrients, calcium, and iron, have a relatively lower MAPE. While all nutrients are important, they noted that potassium and phosphorus, which showed poor performance, are typically less emphasized in general dietary analysis. Experts suggested improving the accuracy of magnesium, vitamin C, and sodium estimates and highlighted potential biases due to dish complexity and portion size estimation. However, higher accuracy was expected in clinical contexts where more precise nutrient intake is essential. In this context, the important nutrients depend on the purpose of conducting a nutritional analysis. For example, potassium and phosphorus are important to kidney patients.
5.4 Study I.4: Comparative Analysis for DietGlance with and without RAG (R2, R3)
@@
5.4.1 Comparing the nutritional analysis results of DietGlance w/ RAG (Study I.3) and w/o RAG
We compared the MAPE of nutrition estimation w/o RAG with the results in Study I.3. As shown in Fig. 8, the system w/ RAG presented lower or close MAPE values in most of the nutrients compared to the system w/o RAG. The system w/o RAG resulted in significantly higher errors in sugar, calcium, and sodium. This indicates that RAG enhances the system’s predictive accuracy, especially for some nutrients that could hidden in sauces, condiments, and bone broth. Since most of the meals collected were Chinese-style, which tend to have higher sodium content due to the frequent use of sodium-rich condiments and seasonings, the lack of the RAG module likely exacerbated the system’s inability to account for the variability in sodium levels. The results demonstrated that while the LLM w/o RAG could achieve satisfactory performance on some nutrients, the incorporation of RAG further enhanced the system’s nutritional analysis capacity, particularly in handling the complexities and variabilities of diverse and culturally sensitive diets.
5.4.2 Comparing dietary suggestion quality w/ and w/o RAG
@@
Metrics for Generative AI. AI-assisted metrics that are widely used for evaluating the AI-generated responses [39] were used to evaluate the quality of dietary suggestions. GPT-4 was used as a grader to rate the responses with a score between 1 and 5, where 1 is the lowest quality and 5 is the highest quality. To compare the responses w/ and w/o RAG, the following metrics from Azure AI Evaluation SDK [^2] Azure AI Evaluation SDK: https://learn.microsoft.com/en-us/azure/ai-studio/how-to/develop/evaluate-sdk were used:
@@
Since suggestions w/ and w/o RAG were evaluated independently, we used GPower [22] to determine the sample size by setting the effect size $d=0.5$ (moderate effect), the significant threshold $\alpha=0.05$, and a statistical power $(1-\beta)=0.8$, resulting in a sample size of 128. Each document was graded twice independently, resulting in 66 grades per group (132 total). We then conducted the Mann-Whitney U Test with a significance level of 0.05 to compare the Relavance-Q, Coherence, and Fluency of dietary suggestions generated w/ and w/o RAG.
@@
6 Study II: Four-Week Longitudinal Study
To evaluate the long-term effect of DietGlance on users’ dietary patterns, we recruited 16 participants from Study I for a 4-week longitudinal study with IRB approval, including 4 females and 12 males (P1, P6, P7, P10, P13, P15, P18, P20, P21, exp., P3, P8, P9, P14, P16, P17, P22, inexp.), aged 24 to 33 ($M=28.3$, $SD=2.4$). Building on the design requirements and findings from Study I, this study aims to explore how exp. and inexp. participants perceive and use DietGlance over time, and how it affects their dietary patterns.
6.1 Study Settings
@@
During the study, participants were encouraged to use Aria Glasses covering as many ingestive behaviors (meals and snacks) as possible to gain comprehensive insights into their habits and use the interface to review their dietary monitoring results. The interface was available 24/7, except from 12:00 am to 5:59 am for potential maintenance and backup, and was tracked via Google Analytics. Meal data older than seven days was archived, with analysis based on the most recent seven days to reflect current dietary patterns. Participants provided weekly feedback through questionnaires and interviews (details in Supplement 4). Participants received about \$17 every week and could review their monitoring and analysis results whenever the interface was available.
We presented the descriptive statistics of the participants’ ratings towards the four design requirements, general usability, user satisfaction, user experience, and nutrient consumption to show the trends across the four weeks. For interviews and textual data, the lead author performed the initial thematic coding and iteratively discussed with other authors to finalize the codes to analyze the impact of DietGlance on participants’ dietary behaviors and their suggestions for improvement.
6.2 Results
@@
Fig. 9: The weekly trends of the scores of exp., inexp., and all participants on (a) Diet identification and logging, (b) Nutritional Analysis, (c) Personalized Dietary Suggestions, and (d) Privacy Protection. The values above the bars are the mean (standard deviation).
6.2.1 Long-Term Perception on Diet Identification and Logging, Nutritional Analysis, Personalized Dietary Suggestions, and Privacy Protection
@@ 6.2.2 Usefulness
6.2.3 Long-Term General Usability, User Satisfaction, and User Experience
Usability improved throughout the study, with SUS scores rising from 76.09 in Study I to 85.00 by the end of Study II, reflecting increased perceived usability. Exp. participants consistently rated the system higher, while inexp. ones showed steady improvement, highlighting adaptability. NPS scores also rose, indicating sustained satisfaction across both groups. User experience, measured through UEQ dimensions, showed positive trends in attractiveness, efficiency, and dependability, while stimulation and novelty declined slightly, reflecting reduced novelty as users became familiar with the system. These findings emphasize the DietGlance’s high usability, satisfaction, and user experience, highlighting the need for tailoring onboarding processes and maintaining engagement, particularly for inexp. users. More details are in Supplement 8.
6.2.4 Long-Term Interaction Behaviors with DietGlance
Participants demonstrated consistent engagement with the system over four weeks, averaging 38.13 usage sessions. Exp. showed slightly higher engagement than inexp. ones. Initial novelty drove higher interactions in Week 1, which gradually stabilized into routine usage by Week 3. The average engagement time per session was 5.95 minutes, with exp. participants spending longer on average, reflecting deeper interaction with the system. The relatively low number of logged updates to diet summaries highlighted the system’s accuracy, while sustained chatbot interactions demonstrated its role in enhancing personalization and feedback. More details are available in the Supplement 9.
@@ 6.2.4 Long-Term Interaction Behaviors with DietGlance
Fig. 10: The consumption trends of different nutrients in Study II. The error bar shows the standard error. Coefficients ($\beta$) and p-values ($p$) from the mixed linear model are shown in the legends with a significant level of 0.05.
6.2.5 Impact of DietGlance on Users’ Dietary Patterns
Nutrient Consumption Trends. To present the nutrition consumption trends across four weeks, we calculated the weekly mean nutrient consumption with min-max normalization. Only nutrients with MAPE < 30% were included. A Mixed Linear Model assessed the trends. As shown in Fig. 10, significant negative trends were observed in energy, total fat, saturated fat, cholesterol, and carbohydrates for both exp. and inexp. groups, suggesting a shift towards a lower-calorie, lower-fat diet. The exp. group also showed a significant reduction in sugar and magnesium, while dietary fiber and iron saw positive trends for both groups, indicating improved consumption of these beneficial nutrients. Fluctuating trends were noted for trans fat, calcium, copper, manganese, and zinc, potentially reflecting variations in food choices and external factors like meal composition and food availability, while the inexp. group showed more variation in sugar and magnesium. Overall, these trends indicate a potential shift towards healthier diets, with reductions in unhealthy fats and sugars, and increases in fiber and iron.
Increasing Understanding of Current Dietary. Both exp. and inexp. participants reported an increased understanding of their dietary habits. P6 (exp.) noticed the high sodium content in cafeteria soup, which aligned with news she had seen about soups being high in salt and purines. She also observed that eating pizza consistently pushed her calorie consumption above the recommended levels. Inexp. participants (e.g., P22) also benefited from understanding their food choices but tended to rely more on general suggestions without deeper dietary analysis. Nearly all participants realized their meals lacked vegetables and fruit, which is a common issue in the location where this study was conducted. P7 (exp.) mentioned that using the system made him more mindful of food nutrient content.
@@ 6.2.5 Impact of DietGlance on Users’ Dietary Patterns
Promoting Healthier Dietary Habits. While many participants had low fruit and vegetable consumption before, both exp. and inexp. participants reported that DietGlance made them intentionally have more vegetables and fruits (P1, P10, P20, exp., P3, P9, P14, P16, P17, P22, inexp.), which also align with the nutrient consumption trends in Fig. 10(f). “Before using this system, I rarely ate vegetables” (P22). Participants also followed system suggestions to achieve dietary goals. For example, P17, aiming to gain weight, increased his consumption of protein-rich foods like tofu and chicken as recommended. P9 noted that she began turning down social meals when not hungry, a change from her previous habit of always joining them regardless of hunger.
@@
6.2.6 Suggestions for Improvement
@@ 7 Discussion
7.1 Implications
7.1.1 Personalization based on Experience, Behaviors, and Environment
@@
The findings emphasize the need to tailor systems to users’ experience levels, interaction behaviors, and environmental contexts. First, exp. participants exhibited higher engagement and satisfaction, emphasizing the need for advanced features like detailed nutritional insights, while simplicity and intuitive interfaces can reduce cognitive load for inexp. users. For example, the system could adapt its workflow and interface to offer advanced functionalities as user experience grows [65].
@@
Second, the system could be enhanced by personalization based on users’ interactions. Image-based food recognition struggles with nuances like cooking methods, seasonings, and visually similar items. Incorporating human-in-the-loop personalization [40] could mitigate this challenge by enabling users to correct misidentified items and share underlying eating habits, enabling the system to learn from these corrections. Additionally, analyzing chatbot conversations can uncover patterns in user preferences and behaviors [60], which, when integrated with clinical data, could further fine-tune LLMs and improve system performance, such as increasing empathy or precision [38].
@@
Moreover, dietary suggestions could further consider environmental factors such as food availability (e.g., university cafeteria menus), financial situations, cultural preferences, and meal timing. As suggested by participants, meal options tailored to commonly visited cafeterias can be incorporated considering the price. Shared meals, common in many cultures such as in Asia, complicate portion size estimation due to variations in individual preferences and dietary needs. While the current system assumes equal portions, adaptive algorithms or informed user input could be incorporated to improve accuracy. Furthermore, conducted with Chinese participants consuming both Western and Eastern cuisines, this study emphasizes the importance of cultural sensitivity in dietary systems. Expanding research to diverse groups would increase the global applicability of such systems.
7.1.2 Balancing Complexity and Relevance in AI-Generative Dietary Suggestions
While the RAG module improves nutrient analysis and dietary suggestions, balancing the complexity of RAG-generated suggestions with their perceived relevance and actionability should be further improved. The complexity of RAG-generated recommendations could be dependent on the complexity of the constructed library. The findings underscore the content curation in building the nutrition library. For example, integrating popular science articles and knowledge summaries targeting the lay public can support the generation of practical, relevant suggestions that align with users’ daily routines and preferences.
@@ 7.1.2 Balancing Complexity and Relevance in AI-Generative Dietary Suggestions
Moreover, participants expressed the need for more detailed context (e.g., the impact of sources) to enhance credibility while avoiding information overload. Future systems should balance source transparency with simplicity. Incorporating techniques like clickable summaries, expandable sections, or collapsible details can help present more in-depth information without overwhelming the user. Future systems can leverage LLMs to offer colloquial explanations of professional terminology [23], making complex nutritional information more accessible to users of varying expertise, thus improving the actionability of suggestions and increasing user engagement.
@@
7.1.3 Enhance the Integration of Multi-Devices and Proactive Intervention

### v3

@@ 5.3.3. Results
5.3.4. Post-Analysis Discussion with Experts
We held a panel discussion with the experts, moderated by the lead researchers, to interpret the quantitative results. The discussion was structured around specific guiding questions focusing on two themes: (1) reasons for low performance in specific nutrients and (2) key nutrients for dietary analysis and suggestions.
@@ 5.3.4. Post-Analysis Discussion with Experts
(2) The experts valued the system’s overall performance for daily dietary tracking, since the important nutrients, such as the macronutrients, calcium, and iron, have a relatively lower MAPE. While all nutrients are important, they noted that potassium and phosphorus, which showed poor performance, are typically less emphasized in general dietary analysis. Experts suggested improving the accuracy of magnesium, vitamin C, and sodium estimates and highlighted potential biases due to dish complexity and portion size estimation. However, higher accuracy was expected in clinical contexts where more precise nutrient intake is essential. In this context, the important nutrients depend on the purpose of conducting a nutritional analysis. For example, potassium and phosphorus are critical to monitor in patients with chronic kidney disease, who often must strictly limit their intake of these minerals.
5.4. Study I.4: Comparative Analysis for DietGlance with and without RAG (R2, R3)
@@
5.4.1. Comparing the nutritional analysis results of DietGlance w/ RAG (Study I.3) and w/o RAG
We compared the MAPE of nutrition estimation w/o RAG with the results in Study I.3. As shown in Fig. 8, the system w/ RAG presented lower or close MAPE values in most of the nutrients compared to the system w/o RAG. The system w/o RAG resulted in significantly higher errors in sugar, calcium, and sodium. This indicates that RAG enhances the system’s predictive accuracy, especially for some nutrients that could be hidden in sauces, condiments, and bone broth. Since most of the meals collected were Chinese-style, which tend to have higher sodium content due to the frequent use of sodium-rich condiments and seasonings, the lack of the RAG module likely exacerbated the system’s inability to account for the variability in sodium levels. The results demonstrated that while the LLM w/o RAG could achieve satisfactory performance on some nutrients, the incorporation of RAG further enhanced the system’s nutritional analysis capacity, particularly in handling the complexities and variabilities of diverse and culturally sensitive diets.
5.4.2. Comparing dietary suggestion quality w/ and w/o RAG
@@
Metrics for Generative AI. AI-assisted metrics that are widely used for evaluating the AI-generated responses (Li et al., 2024b) were used to evaluate the quality of dietary suggestions. GPT-4 was used as a grader to rate the responses with a score between 1 and 5, where 1 is the lowest quality and 5 is the highest quality. To compare the responses w/ and w/o RAG, the following metrics from Azure AI Evaluation SDK (Microsoft, ) were used:
@@
Since suggestions w/ and w/o RAG were evaluated independently, we used GPower (Faul et al., 2009) to determine the sample size by setting the effect size $d=0.5$ (moderate effect), the significant threshold $\alpha=0.05$, and a statistical power $(1-\beta)=0.8$, resulting in a sample size of 128. Each document was graded twice independently, resulting in 66 grades per group (132 total). We then conducted the Mann-Whitney U Test with a significance level of 0.05 to compare the Relavance-Q, Coherence, and Fluency of dietary suggestions generated w/ and w/o RAG.
@@
In summary, this comparative analysis demonstrates the specific and critical advantages of integrating the RAG module. While the baseline LLM (w/o RAG) could generate plausible responses, the RAG-grounded system showed significantly better performance in areas essential for a health application. Specifically, for nutritional analysis, RAG provided more accurate estimations for key nutrients like sugar, calcium, and sodium. For dietary suggestions, the RAG-enhanced version received significantly higher expert ratings for Accuracy. Furthermore, unlike the baseline, the RAG module provides references for its claims, ensuring that the system’s advice is traceable and verifiable, which is a critical enhancement for preventing misinformation in health-sensitive applications.
6. Study II: Four-Week Longitudinal Study
To evaluate the long-term effect of DietGlance on users’ dietary patterns, we recruited 16 participants from Study I in Hong Kong for a 4-week longitudinal study with IRB approval, including 4 females and 12 males (P1, P6, P7, P10, P13, P15, P18, P20, P21, exp., P3, P8, P9, P14, P16, P17, P22, inexp.), aged 24 to 33 ($M=28.3$, $SD=2.4$). Building on the design requirements and findings from Study I, this study aims to explore how exp. and inexp. participants perceive and use DietGlance over time, and how it affects their dietary patterns.
6.1. Study Settings
@@
During the study, participants were encouraged to use Aria Glasses covering as many ingestive behaviors (meals and snacks) as possible to gain comprehensive insights into their habits and use the interface to review their dietary monitoring results. The interface was available 24/7, except from 12:00 am to 5:59 am for potential maintenance and backup, and was tracked via Google Analytics. Meal data older than seven days was archived, with analysis based on the most recent seven days to reflect current dietary patterns. Participants provided weekly feedback through questionnaires and interviews (details in Appendix E). Participants received about \$17 every week and could review their monitoring and analysis results whenever the interface was available.
We presented the descriptive statistics of the participants’ ratings towards the four design requirements, general usability, user satisfaction, user experience, and nutrient consumption, to show the trends across the four weeks. For interviews and textual data, the lead author performed the initial thematic coding and iteratively discussed with other authors to finalize the codes to analyze the impact of DietGlance on participants’ dietary behaviors and their suggestions for improvement.
6.2. Results
@@
Fig. 9. The weekly trends of the scores of exp., inexp., and all participants on (a) Diet identification and logging, (b) Nutritional Analysis, (c) Personalized Dietary Suggestions, and (d) Privacy Protection. The values above the bars are the mean (standard deviation).
6.2.1. Long-Term Perception on Diet Identification and Logging, Nutritional Analysis, Personalized Dietary Suggestions, and Privacy Protection
@@ 6.2.2. Usefulness
6.2.3. Long-Term General Usability, User Satisfaction, and User Experience
Usability improved throughout the study, with SUS scores rising from 76.09 in Study I to 85.00 by the end of Study II, reflecting increased perceived usability. Exp. participants consistently rated the system higher, while inexp. ones showed steady improvement, highlighting adaptability. NPS scores also rose, indicating sustained satisfaction across both groups. User experience, measured through UEQ dimensions, showed positive trends in attractiveness, efficiency, and dependability, while stimulation and novelty declined slightly, reflecting reduced novelty as users became familiar with the system. These findings emphasize the DietGlance’s high usability, satisfaction, and user experience, highlighting the need for tailoring onboarding processes and maintaining engagement, particularly for inexp. users. More details are in Appendix I.
6.2.4. Long-Term Interaction Behaviors with DietGlance
Participants demonstrated consistent engagement with the system over four weeks, averaging 38.13 usage sessions. Exp. showed slightly higher engagement than inexp. ones. Initial novelty drove higher interactions in Week 1, which gradually stabilized into routine usage by Week 3. The average engagement time per session was 5.95 minutes, with exp. participants spending longer on average, reflecting deeper interaction with the system. The relatively low number of logged updates to diet summaries highlighted the system’s accuracy, while sustained chatbot interactions demonstrated its role in enhancing personalization and feedback. More details are available in the Appendix J.
@@ 6.2.4. Long-Term Interaction Behaviors with DietGlance
Fig. 10. The consumption trends of different nutrients in Study II. The error bar shows the standard error. Coefficients ($\beta$) and p-values ($p$) from the mixed linear model are shown in the legends with a significant level of 0.05.
6.2.5. Impact of DietGlance on Users’ Dietary Patterns
Nutrient Consumption Trends. To present the nutrition consumption trends across four weeks, we calculated the weekly mean nutrient consumption with min-max normalization. Only nutrients with MAPE ¡ 30% were included. A Mixed Linear Model assessed the trends. As shown in Fig. 10, significant negative trends were observed in energy, total fat, saturated fat, cholesterol, and carbohydrates for both exp. and inexp. groups, suggesting a shift towards a lower-calorie, lower-fat diet. The exp. group also showed a significant reduction in sugar and magnesium, while dietary fiber and iron saw positive trends for both groups, indicating improved consumption of these beneficial nutrients. Fluctuating trends were noted for trans fat, calcium, copper, manganese, and zinc, potentially reflecting variations in food choices and external factors like meal composition and food availability, while the inexp. group showed more variation in sugar and magnesium. Overall, these trends indicate a potential shift towards healthier diets, with reductions in unhealthy fats and sugars, and increases in fiber and iron.
Increasing Understanding of Current Dietary. Both exp. and inexp. participants reported an increased understanding of their dietary habits. P6 (exp.) noticed the high sodium content in cafeteria soup, which aligned with news she had seen about soups being high in salt and purines. She also observed that eating pizza consistently pushed her calorie consumption above the recommended levels. Inexp. participants (e.g., P22) also benefited from understanding their food choices but tended to rely more on general suggestions without deeper dietary analysis. Nearly all participants realized their meals lacked vegetables and fruit, which is a common issue in Hong Kong, where this study was conducted. P7 (exp.) mentioned that using the system made him more mindful of food nutrient content.
@@ 6.2.5. Impact of DietGlance on Users’ Dietary Patterns
Promoting Healthier Dietary Habits. While many participants had low fruit and vegetable consumption before, both exp. and inexp. participants reported that DietGlance made them intentionally have more vegetables and fruits (P1, P10, P20, exp., P3, P9, P14, P16, P17, P22, inexp.), which also aligns with the nutrient consumption trends in Fig. 10(f). “Before using this system, I rarely ate vegetables” (P22). Participants also followed system suggestions to achieve dietary goals. For example, P17, aiming to gain weight, increased his consumption of protein-rich foods like tofu and chicken as recommended. P9 noted that she began turning down social meals when not hungry, a change from her previous habit of always joining them regardless of hunger.
@@
6.2.6. Suggestions for Improvement
Based on participant feedback, we identified five key areas for improvement:
@@ 7. Discussion
7.1. Implications
7.1.1. Personalization based on Experience, Behaviors, and Environment
@@
The findings emphasize the need to tailor systems to users’ experience levels, interaction behaviors, and environmental contexts. First, exp. participants exhibited higher engagement and satisfaction, emphasizing the need for advanced features like detailed nutritional insights, while simplicity and intuitive interfaces can reduce cognitive load for inexp. users. For example, the system could adapt its workflow and interface to offer advanced functionalities as user experience grows (Saif et al., 2024).
@@
Second, the system could be enhanced by personalization based on users’ interactions. Image-based food recognition struggles with nuances like cooking methods, seasonings, and visually similar items. Incorporating human-in-the-loop personalization (Lin et al., 2020) could mitigate this challenge by leveraging the user’s existing ability to correct misidentified items to dynamically fine-tune the model. This would allow the system to learn from these corrections over time, reducing the need for repeated manual interventions. Additionally, analyzing chatbot conversations can uncover patterns in user preferences and behaviors (Qin et al., 2024), which, when integrated with clinical data, could further fine-tune LLMs and improve system performance, such as increasing empathy or precision (Li et al., 2024a; Jiang et al., 2026).
@@
Moreover, dietary suggestions could further consider environmental factors such as food availability (e.g., university cafeteria menus), financial situations, cultural preferences, and meal timing. As suggested by participants, meal options tailored to commonly visited cafeterias can be incorporated considering the price. Shared meals, common in many cultures such as in Asia, complicate portion size estimation due to variations in individual preferences and dietary needs. While the current system assumes equal portions, adaptive algorithms or informed user input could be incorporated to improve accuracy. Furthermore, while our participants consumed diverse cuisines, their shared cultural background (Chinese) limits the generalizability of the observed behavioral patterns and preferences (e.g., regarding shared meals). Therefore, expanding research to more diverse cultural groups is a critical next step to increase the global applicability of such systems.
7.1.2. Balancing Complexity and Relevance in AI-Generative Dietary Suggestions
While the RAG module improves nutrient analysis and dietary suggestions, balancing the complexity of RAG-generated suggestions with their perceived relevance and actionability should be further improved. The complexity of RAG-generated recommendations could be dependent on the complexity of the constructed library. This highlights a key limitation of RAG: its output is highly dependent on the quality, complexity, and relevance of the retrieved documents. Our expert evaluation revealed that our library’s inclusion of academic articles, while ensuring accuracy, produced recommendations that were at times overly complex, thereby adversely affecting their perceived relevance and actionability. The curation of the knowledge base is therefore a critical, non-trivial step that impacts system performance. For example, integrating carefully vetted popular science articles and knowledge summaries (e.g., from reputable health organizations) targeting the lay public can support the generation of practical, relevant suggestions that align with users’ daily routines and preferences. However, to mitigate the risk of misinformation, such content must undergo rigorous verification against clinical guidelines and be presented with clear user disclaimers. Complementing this, future work could also leverage Chain-of-Thought prompting (Wei et al., 2022) and style transfer techniques (Toshevska and Gievska, 2025) to decompose dense clinical guidelines into digestible formats, automatically translating technical jargon into plain language for the general audience.
@@ 7.1.2. Balancing Complexity and Relevance in AI-Generative Dietary Suggestions
Moreover, participants expressed the need for more detailed context (e.g., the impact of sources) to enhance credibility while avoiding information overload. Future systems should balance source transparency with simplicity. Incorporating techniques like clickable summaries, expandable sections, or collapsible details can help present more in-depth information without overwhelming the user. Future systems can leverage LLMs to offer colloquial explanations of professional terminology (Fok et al., 2024), making complex nutritional information more accessible to users of varying expertise, thus improving the actionability of suggestions and increasing user engagement.
Furthermore, as an alternative to RAG, fine-tuning an LLM on a high-quality, domain-specific dataset may present a certain degree of ability to internalize general nutritional knowledge and adopt a more consistent, practical tone. However, fine-tuning alone shares key limitations with baseline LLMs. First, it demands ongoing maintenance; unlike RAG, knowledge cannot be easily updated without retraining the model. Second, fine-tuning requires a large volume of high-quality training data (e.g., expert-verified QA pairs), which is labor-intensive to curate. Finally, fine-tuned outputs remain unverifiable, as they usually cannot be explicitly traced back to a specific trusted source. A promising direction for future work could be an agentic framework. This would involve an agent that could intelligently route user queries, deciding when to retrieve verifiable facts via RAG from the latest resources (e.g., for specific nutrient data) and when to rely on its fine-tuned internal knowledge (e.g., for general encouragement or stylistic responses). This could offer a more flexible solution to balance accuracy and actionability.
@@
7.1.3. Enhance the Integration of Multi-Devices and Proactive Intervention